OpenStudy (anonymous):

7(3x+9) = 11 - (x+3) x= -5/2 is this correct simplified?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks!

OpenStudy (turingtest):

21x+63=11-x-3 22x=-55 x=-55/22=-5/2 looks like it